OpenStudy (anonymous):

how do you do ^3radical 8?

OpenStudy (anonymous):

\[6\sqrt{2}\]

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

\[3\sqrt{8} = 3\sqrt{2^{2}\times2} = 3\times 2 \sqrt{2} = 6\sqrt{2}\]

OpenStudy (anonymous):

where as if you're asking about this thn the \[\sqrt[3]{8} = 2\]